OpenStudy (anonymous):

Limit as x approaches infinity of 8x[ln(x+9) - ln(x)]

OpenStudy (anonymous):

How can I rewrite it as a fraction in order to apply L'Hopital's rule?

hartnn (hartnn):

you can bring x in the denominator and write it as 1/x

hartnn (hartnn):



OpenStudy (anonymous):

Right but then plugging in infinity I would get infinity/0 right? And thats not defined

hartnn (hartnn):

you wanted to apply LH right ? so first differentiate numerator and denominator separately later think about plugging it, after some simplification

OpenStudy (anonymous):

Thats true but in order to use LH we have to first make sure that we have an indeterminate form and oo/0 isn't it

hartnn (hartnn):

right, did you try the substitution x =1/y ? when x-> infinity, y->0

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty }\left(\ln(x+9)-\ln(x)\right)}\) \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~\infty }\left(\ln\frac{x+9}{x}\right)}\) \(\large\color{slate}{\displaystyle\left(\ln\left(\lim_{x \rightarrow ~\infty }\frac{x+9}{x}\right)\right)}\)

OpenStudy (solomonzelman):

so you are taking the ln of that limit

OpenStudy (solomonzelman):

0

OpenStudy (solomonzelman):

oh 8x !!1soprry

OpenStudy (anonymous):

Yes thats right! Oh but where's the 8x?

OpenStudy (solomonzelman):

yes yes, my bad

OpenStudy (anonymous):

And so is that also one of the limit properties? Exchanging the places of ln and limit?

OpenStudy (solomonzelman):

yes limit properties allow thatr

OpenStudy (anonymous):

Alright so using the same idea where would the 8x go?

hartnn (hartnn):

its not exchanging... Limit can be dragged inside of only those functions which are continuous

hartnn (hartnn):

can you give the substitution one shot? the substitution x =1/y ? when x-> infinity, y->0